rk Question 19 of 20 Check My Work (3 remaining) here to read the eBook: Net Present Value (NPV) here to read the eBook: Internal Rate of Return (IRR) AND IRR store has 5 years remaining on its lease in a mall. Rent is $1,900 per month, 60 payments remain, and the next payment is due in 1 month. The malls owner plans to sell the erty in a year and wants rent at that time to be high so that the property will appear more valuable. Therefore, the store has been offered a great deal (owners words) on a nem ear ise, the ne lease calls for n rent for 9 months, then payments of $2.600 per month for the next 51 months. The lease cannot be broken, and the stores wACC is 12% or per roonth) should the new leage be accepted? (Hint: Be sure to use 1% per month.) f the store owner decided to bargain with the malls owner over the new lease payment, what new lease payment would make the store owner indifferent between the new and old leases ? (Hint: Find FV of the old leases original cost at t-9: then treat this as the PV of a 51-period annuity whose payments represent the rent during months 10 to 60.) Round r answer to the nearest cent. Do not round your intermediate calculations c. The store owner is not sure of the 12% WACC-it could be higher or lower. At what nominal WAcc would the store owner be indifferent between the two leases?( differences between the two payment streams; then find its IRR) Round your answer to two decimal places. Do not round your intermediate calculations DOLL
Answer #1

Let's calculate the present value of all these leases using PV function

N = 60, I/Y = 1%, PMT = 1,900, FV = 0 => Compute PV = $85,414.57 is the original PV

N = 51, I/Y = 1%, PMT = 2,600, FV = 0 => Compute PV = $103,475.15 is the PV after 9 months

PV today = 103,475.15 / (1 + 1%)^9 = $94,611.45

The new lease is more expensive to you and hence, you should not accept the offer.

b) PV of original lease in 9 months = 85,414.57 x (1 + 1%)^9 = $93,416.66

Indifference lease can be calculated using PMT function

N = 51, I/Y = 1%, PV = 93,416.66, FV = 0 => Compute PMT = $2,347.26 should be the indifference lease.

c) Using trial and error method or excel solver or IRR function, we can calculate the indifference rate, which is equal to 32.91% at which PV for both lease are equal.